The length of the rectangle is three times more than its width.Find its length and width when its area is 270 cm^
Answer by Akemi(2)

First, represent the variables...
l is for length ; w is for width
l = 3+w
w = w
how to solve it :
(3+w)(w) = 270
w^2 + 3w = 270
w^2 + 3w - 270 =0
(w+18)(w-15)
answer:
l = 3+w(15) = 3+15 = 18
w= 15
therefore,
the length is 18 and width is 15.
:D